Suppose he production function takes the form Y : F KN where Y he quantity of ou put Kis the quanti o capital goods and is quantity of a Suppose specifically that Recall that the marginal product of labor is the the extra amount of output that results from using one extra unit of labor, while holding constant the amount of capital used. Suppose N-9 and K-16. Then, to two decimal places, the marginal product of labor is K N Suppose instead that N-9 and K-25. Then, to two decimal places, the marginal product of labor is
Answer #1

Marginal product of labour = MPN = dF/dN = 5(K/N)0.5

When N = 9 and K = 16

MPN = 5 x 4/3 = 6.67

When N = 9 and K = 25

MPN = 5 x 5/3 = 8.33